Solve by completing the square. x2 8x + 11 = 0
Answer #1

a² Sa+ll=0 ær -sat16-16711=0 8 a .2.0.4+42 -5= Ca²q-204+48) -5=0 (a-uge_5=0 (a-478 = g La-u) tra a=4 I to a=4-15, 4+05